An analyst believes that economic conditions during the next year will either be strong, normal, or weak, and she thinks that the Corrigan Company's returns will have the following probability distribution.

Conditions Probability (%) Return (%)

Strong 30 30

Normal 40 15

Weak 30 -10


What is Corrigan’s standard deviation of returns?


1
Expert's answer
2022-03-27T18:52:15-0400

The average returns

"R =\frac{(30+15-10)}{3} = 11.667"%


Finding the varaince;

Variance"= \sum" "\frac {(Returns - R)^2}{n-1}"

"= \frac{(30-11.667)^2+(15-11.667)^2 + (-10-11.667)^2}{3-1} = 408.33"

Finally, from the above variance we can compute the standard deviation as;

"SD =" "\sqrt{ Variance}"

"= \sqrt{408.33}"

"= 20.21"%